Why belong exponential family for poisson distribution or geometric distribution?
Why belong exponential family for poisson distribution
What is the probability of pulling a queen out of a shuffled card deck?
There is a 1 (one) in 13 (thirteen) chance that you will pull out a queen from a deck of 52 (fifty-two) cards.
2/36 what you rolled the first time has nothing to do with the next roll. It's an independent event.
3/12*3/11 = 9/132, or 6.818%.
It is 0.4
Why are important in random variables?
It might help if you specified why WHAT was important in random variables.
Does something that can never happen have a probability of 0?
Yes, a probability of 0 is something that cannot happen.
When rolling two six sided number cubes the probability that on one roll you get at least a nine?
It is 5/18.
What is the definition of probability?
The likelihood that a particular event will occur is called probability
The ratio of the wanted outcomes divided by the possible outcomes gives a fraction or ratio usually expressed as a percentage.
The likelihood of an event occuring.
The amount of time that a certain type of battery functions is a random variable?
A six-sided die, in which each side is equally likely to appear, is repeatedly rolled until the total of all rolls exceeds 400. Approximate the probability that this will require more than 140 rolls.
Number of ways 5 objects can be arranged in a row?
5! is the symbol for Factorial 5 , ie 5 x 4 x 3 x 2 = 120 different ways
What is an example of probability?
If you have 10 marbles in a bag, 5 red 2 blue, and 3 purple, and you randomly draw one out the percentage of it being Red=5/10=1/2 Blue=2/10=1/5 Purple=3/10
The answer to the question, as stated, is two. You are drawing from the same regular deck. If the questioner meant same suit, the answer is 5. The first four could be one of each of the four suits but the fifth must belong to one of the suits you have already drawn.
What is the probability of rolling the sum of 3 with 2 dice?
1:18. There are 36 potentials and only two can add to 3. Thus one in eighteen
What is the probability of getting a red card and a red card assuming the red card is replaced?
The probability of drawing two red cards with replacement is (1 in 2)2 or (1 in 4) or 0.25
Define statistical inference. Give one example of statistical inference?
Define statistical inference and give an example